### v1

# Fluctuations of the free energy of the spherical Sherrington–Kirkpatrick model with heavy-tailed interaction

Abstract
We consider the 2-spin spherical Sherrington–Kirkpatrick model where the interactions between the spins are given as random variables with heavy-tailed distribution. We prove that the free energy exhibits sharp phase transition, depending on the location of the largest eigenvalue of the interaction matrix. We also prove the order of the limiting free energy and the limiting distribution of the fluctuation of the free energy for both regimes.
1 Introduction
The spin glass model, initially introduced in the seminal work of Edwards and Anderson [12], has evolved significantly over time. One notable advancement was the development of its mean-field variant, the Sherrington–Kirkpatrick (SK) model [23]. Despite its apparent simplicity, the SK model captures many fundamental properties of the real-world spin glass model, which has led to its widespread study and application in a variety of contexts. A critical aspect of the SK model is its free energy, which can be conceptualized through the replica symmetry breaking method, notably the Parisi formula [22], which was later mathematically validated [14, 26].
A notable variant of the SK model is the spherical SK model, introduced by Kosterlitz, Thouless, and Jones [17]. It replaces the standard spin vectors with those uniformly distributed on a sphere. The spherical SK model has garnered attention for its applications in statistical mechanics and other fields like information theory [21], with significant progress made in understanding its mathematical properties, including its limiting free energy [25] and the fluctuation of the free energy [5]. For other variants of the SSK model, we refer to [4] and references therein.
While the exact formulas for the limit and the fluctuation of the free energy may depend on the detail of the model, we can find one common feature in these models: The limiting free energy is a continuous function of the inverse temperature. Technically, it can be understood as follows: (1) by the integral representation formula (e.g., Lemma 1.3 in [5]), the partition function depends only on the eigenvalues of the interaction matrix, (2) the eigenvalues are close to the deterministic locations determined by their limiting empirical spectral distribution (ESD), (3) as a result, the limiting free energy can be well-approximated by a deterministic function of the inverse temperature.
In this paper, we consider a variant of the SSK model where the interactions between the spins are given as random variables with heavy-tailed distribution and investigate how this heavy-tailed interaction can alter the continuity of the limiting free energy. Indeed, there are several reasons that we can believe that the heavy-tailed SSK model can exhibit a feature that is significantly different from other SSK models. First, the ESDs of the heavy-tailed random matrices are markedly different from corresponding ‘light-tailed’ Wigner random matrices. Second, recent advancements have been made in calculating the free energy fluctuation for the SK model with heavy-tailed interactions [10], revealing significant differences from the usual SK model.
Before describing our model more in detail, we recall the key results for the usual SSK model [5]. The Hamiltonian for the SSK model is defined by
$H_{N}(\sigma)=-\frac{1}{\sqrt{N}}\sum_{i,j=1}^{N}H_{ij}\sigma_{i}\sigma_{j}=-\frac{1}{\sqrt{N}}\langle\sigma,H\sigma\rangle$ (1.1)
for a real symmetric random matrix $H$, where $\sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{N})\in S_{N}:=\{x\in\mathbb{R}^{N}:x_{1}^{2}+\dots+x_{N}^{2}=N\}$. The entries of $H$ are with mean 00 and variance $1$, so that the largest eigenvalue of $H$ is of order $\sqrt{N}$ with high probability. The partition function and the free energy of the SSK model is defined as
$F_{N}\equiv F_{N}(\beta)=\frac{1}{N}\log{Z_{N}},\quad Z_{N}=\int_{S_{N}}e^{\beta H_{n}(\sigma)}\mathrm{d}\sigma,$ (1.2)
where $\beta$ is the inverse temperature and $\mathrm{d}\sigma$ is the Haar measure (normalized uniform measure) on $S_{N}$. It is known that
$F_{N}(\beta)\to F(\beta)\equiv\begin{cases}\beta^{2}&\text{ if }0<\beta\leq 1/2,\\ 2\beta-\frac{\log(2\beta)+\frac{3}{2}}{2}&\text{ if }\beta>1/2,\end{cases}$ (1.3)
as $N\to\infty$. Moreover, the fluctuation of the free energy is given by Gaussian law of order $N^{-1}$ when $\beta<1/2$ and GOE Tracy–Widom law of order $N^{-2/3}$ when $\beta>1/2$.
In this paper, we further assume that the entries $H$ are heavy-tailed random variables in the sense that $\mathbb{P}(|H_{ij}|>u)\sim u^{-\alpha}$ for some $\alpha\in(0,2)$. The exponent $\alpha$ determines the heaviness of the tail of the distributions. With heavy-tailed entries, unlike the light-tailed model, the largest eigenvalue is not of order $1$ but grows with $N$. We thus need to introduce an additional normalization factor $b_{N}$, which is order asymptotically $N^{2/\alpha}$, and consider the partition function defined as
$Z_{N}=\int_{S_{N}}e^{b_{N}^{-1}\beta H_{N}(\sigma)}\mathrm{d}\sigma.$ (1.4)
See Section 2.1, especially Definition 2.5, for the precise definition of the model and the partition function $Z_{N}$ with the normalization factor $b_{N}$. We remark that the normalization factor $b_{N}$ ensures that the extreme eigenvalues of $b_{N}^{-1}H$ is of order $1$; see [3, 24] for more detail.
1.1 Main contribution
In this paper, we prove the behavior of the (log-) partition function in the “high temperature regime” and the “low temperature regime”. We find that
  • (Theorem 2.6.(ii)) In the high temperature regime, $\log Z_{N}$ converges in distribution to a random variable, independent of $N$, as $N\to\infty$.
  • (Theorem 2.6.(iii)) In the low temperature regime, $N^{-1}\log Z_{N}$ converges in distribution to another random variable, independent of $N$, as $N\to\infty$.
When compared to the behavior of the free energy for the usual SSK model, we notice several contrasts. First, while for the usual SSK model the size of $\log Z_{N}$ is of order $N$ regardless of the temperature, it is markedly different for our model in the high temperature regime and the low temperature regime - the former is of order $1$ whereas the latter is of order $N$. Second, for the usual SSK model the fluctuation of $\log Z_{N}$ is much smaller than its deterministic limit, they are of comparable order for our model.
The transition between the high temperature regime and the low temperature regime is even more striking for our model. For the usual SSK model, the behavior of the free energy is solely determined by the inverse temperature. It is in the high temperature regime if $\beta$ is below a certain threshold, $1/2$ in (1.3), and in the low temperature regime if $\beta$ is below the threshold; no randomness is involved in the phase transition. On the other hand, with heavy-tailed interaction matrix, the distinction between the two phases are more subtle. In Theorem 2.6, we prove that the high temperature regime is when the largest eigenvalue of the interaction matrix is above a certain threshold determined by $\beta$ and the low temperature regime is when the largest eigenvalue is below the threshold. It in particular implies that the size of $\log Z_{N}$ is of order $1$ with probability $P_{\beta}$ and of order $N$ with probability $1-P_{\beta}$. The change of the inverse temperature only changes the probability $P_{\beta}$.
For the proof of the main result, we adapt the strategy introduced in [5]. In this approach, the partition function is first written as a complex integral involving the eigenvalues of the interaction matrix (see (2.4)). Then, with the aid of several estimates on the eigenvalues of heavy-tailed matrices, the integral is approximated by applying the method of steepest descent. In the high temperature regime, the (log-) partition function can be approximated by a linear spectral statistics (LSS) of the eigenvalues, while it is governed by the largest eigenvalue in the low temperature regime.
The main technical difficulty in the analysis is due to that some spectral properties of heavy-tailed matrices are not known, e.g., its LSS with respect to a logarithmic function. To handle it, we notice that the fluctuation of several quantities such as the trace and the Hilbert-Schmidt norm of a heavy-tailed matrix is governed by a few extreme entries of the matrix. Then, by applying well-known facts about heavy-tailed distributions, we can analyze the LSS with respect to a logarithmic function that is essential in the analysis for high temperature regime.
1.2 Related works
The SSK model was first introduced by Kosterlitz, Thouless, and Jones [17]. The limiting free energy for the SSK model was computed (non-rigorously) in [17] and proved by Talagrand [25]. The fluctuation of the free energy for the SSK model was proved in [5]. Corresponding results for several variants of the SSK model were proved in [5, 6, 7, 20, 16]. Several other results were also proved, including the near-critical behavior [8, 18, 15] and the overlap [19]. For more results on the SSK model and its variants, we refer to [4] and references therein.
The spectral properties of the heavy-tailed random matrices have been extensively studied in random matrix theory. The behavior of the largest eigenvalue was first proved by Soshnikov [24] for the case $\alpha<2$ and later generalized to the case $\alpha<4$ by Auffinger, Ben Arous, and Péché [3]. Several other results were proved, including the spectrum normalization [2], a central limit theorem for the Stieltjes transform [9], and bulk universality [1].
1.3 Organization of the paper
In Section 2, we define our model and state the main results. In section 3, we introduce several important properties for heavy-tailed random matrices. In Sections 4 and 5, we prove the main results in the high temperature regime and the low temperature regime, respectively. Several technical details, including the proofs of the lemmas, can be found in Appendix.
2 Definitions and main results
In this section, we precisely define our model and state the main results.
Remark 2.1 (Notational remark).
Throughout the paper, we use the standard asymptotic notations such as $O$, $o$, and $\Theta$ as $N\to\infty$. We denote by the symbol $\Rightarrow$ the convergence in distribution as $N\to\infty$. We use the shorthand notation $\sum_{i}=\sum_{i=1}^{N}$.
2.1 Definitions
We begin by defining the heavy-tailed random variables.
Definition 2.2 (Heavy-tailed random variable).
We say that a random variable $X$ is heavy-tailed with exponent $\alpha$ if $\mathbb{P}(|X|>u)=L(u){u^{-\alpha}}$ for some $\alpha\in(0,1)$ and some slowly varying function $L$, i.e. $\lim_{x\to\infty}L(tx)/L(x)=1$ for all $t$. We further assume that for any $\delta>0$, there exists $x_{0}$ such that $L(x)e^{x^{\delta}}$ is an increasing function on $(x_{0},\infty)$.
Remark 2.3.
In Definition 2.2, the assumption on the monotonicity of the function $L(x)e^{x^{\delta}}$ is purely technical, which guarantees that the difference $\lambda_{1}-\lambda_{2}$ is bounded below with high probability (in Lemma 3.6). This assumption is satisfied by many slowly varying functions considered in the literature, especially poly-log functions $(\log(x))^{p}$ for some $p\in\mathbb{R}$.
Note that a heavy-tailed random variable with exponent $\alpha$ has a finite $(\alpha-\epsilon)$ moment for any $\epsilon\in(0,\alpha)$. We next define a Wigner matrix with heavy-tailed entries.
Definition 2.4 (Heavy-tailed Wigner matrix).
We say that an $N\times N$ real symmetric matrix $M$ is a heavy-tailed Wigner matrix with exponent $\alpha$ if its upper triangular entries $M_{ij}(1\leq i\leq j\leq N)$ are independent and identically distributed (i.i.d.) heavy-tailed random variables with exponent $\alpha$, defined in Definition 2.2.
Our model is spherical SK model in which the interactions between the spins are i.i.d. heavy-tailed random variables. Its partition function and the free energy are defined as follows:
Definition 2.5 (Partition function).
For an $N\times N$ heavy-tailed Wigner matrix $M$ with exponent $\alpha$, let $b_{N}$ be a normalization factor defined by
$b_{N}=\inf\{t:\mathbb{P}(|M_{11}|>t)\leq\frac{2}{N(N+1)}\}$ (2.1)
where $L$ is the slowly varying function associated the entries of $M$ as in Definition 2.2. We define the partition function associated with $M$ at inverse temperature $\beta$ by
$Z_{N}\equiv Z_{N}(\beta)=\int_{S_{N}}\exp(\beta b_{N}^{-1}\langle x,Mx\rangle)\,\mathrm{d}\Omega_{N}(x),$ (2.2)
where $S_{N}:=\{x\in\mathbb{R}^{N}:x_{1}^{2}+\dots+x_{N}^{2}=N\}$ and we denote by $\mathrm{d}\Omega_{N}$ the Haar measure (normalized uniform measure) on $S_{N}$.
2.2 Main results
Our main result is the following theorem on the limiting behavior of the free energy.
Theorem 2.6.
Suppose that $M$ is an $N\times N$ heavy-tailed Wigner matrix with exponent $\alpha$ as in Definition 2.4. Then, there exist events $F_{1}$ and $F_{2}$ such that
  1. (i)
    Events $F_{1}$ and $F_{2}$ are mutually exclusive with
    $\lim_{N\to\infty}\mathbb{P}(F_{1})=\exp(-(2\beta)^{\alpha}),\quad\lim_{N\to\infty}\mathbb{P}(F_{2})=1-\exp(-(2\beta)^{\alpha})$
  2. (ii)
    Given $F_{1}$, for the partition function $Z_{N}$ in Definition 2.5, as $N\to\infty$
    $(\log Z_{N})|F_{1}\Rightarrow T$
    for some ($N$-independent) random variable $T$.
  3. (iii)
    Given $F_{2}$, for the partition function $Z_{N}$ in Definition 2.5, as $N\to\infty$
    $(\frac{1}{N}\log Z_{N})|F_{2}\Rightarrow\beta X-\frac{1}{2}\log(2e\beta X)$
    for some ($N$-independent) random variable $X$.
The precise definitions of $F_{1}$ and $F_{2}$ in Theorem 2.6 are
$F_{1}=\{\lambda_{1}<\frac{b_{N}}{2\beta}\},\quad F_{2}=\{\lambda_{1}>\frac{b_{N}}{2\beta}\}.$ (2.3)
As we discussed in Introduction, the events $F_{1}$ and $F_{2}$ correspond to the high temperature regime and the low temperature regime, respectively. We remark that $X$ in the low temperature regime is a random variable such that $X>1/(2\beta)$ and its distribution is given by
$\mathbb{P}(X>u)=\exp(-x^{-\alpha}+(2\beta)^{\alpha}).$
We also remark that the random variable $T$ in the high temperature regime has undefined mean.
Proof of Theorem 2.6.
See Proposition 4.1 in Section 4 and Proposition 5.1 in Section 5. ∎
2.3 Strategy of the proof
For the proof of Theorem 2.6, we adapt the integral representation formula for the partition function, introduced in [17] and proved in [5]. Let $\lambda_{1}\geq\lambda_{2}\geq\dots\lambda_{N}$ be the eigenvalues of $M$. In Lemma 3.16, we prove that
$Z_{N}=C_{N}\frac{1}{i}\int_{\gamma-i\infty}^{\gamma+i\infty}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z$ (2.4)
for any $\gamma>\lambda_{1}$, where
$G(z)=2\beta z-\frac{b_{N}}{N}\sum_{i}\log(z-\lambda_{i}),\quad C_{N}=\frac{\Gamma(N/2)b_{N}^{N/2-1}}{2\pi(N\beta)^{N/2-1}}.$ (2.5)
To analyze the asymptotic behavior of the right side of (2.4), in the viewpoint of the method of steepest descent, the best candidate for $\gamma$ would be the critical point of the function $G$, i.e., the solution of the equation
$G^{\prime}(z)=2\beta-\frac{b_{N}}{N}\sum_{i}\frac{1}{z-\lambda_{i}}=0.$ (2.6)
Note that $G^{\prime}(z)$ is an increasing function of $z$ and (2.6) has a unique solution in $(\lambda_{1},\infty)$.
While $G^{\prime}(z)$ is a random function, since $\lambda_{i}=O(1)$ for most of the eigenvalues $\lambda_{i}$, it is possible to approximate $G^{\prime}$ by a deterministic function with negligible error. We can then approximate the solution of the equation (2.6) by a deterministic number $b_{N}/(2\beta)$. Thus, if $\lambda_{1}<b_{N}/(2\beta)$, which we call the high temperature regime, we can analyze the asymptotic behavior of $Z_{N}$ by choosing $\gamma=b_{N}/(2\beta)$. In Section 4, we consider this case and prove the result on the fluctuation of $\log Z_{N}$, Proposition 4.1, by considering the linear spectral statistics of $M$ involving the function $G(z)$ in (2.5).
If $\lambda_{1}>b_{N}/(2\beta)$, on the other hand, we cannot choose $\gamma$ to be $b_{N}/(2\beta)$. In this case, which we call the low temperature regime, we prove that the solution of the equation (2.6) sticks to the largest eigenvalue $\lambda_{1}$, and the behavior of $Z_{N}$ is directly related to that of $\lambda_{1}$ with negligible error. In Section 5, we prove the limiting behavior of $Z_{N}$ as in Proposition 4.1, by considering the behavior of $\lambda_{1}$.
For both the high temperature regime and the low temperature regime, it is crucial in the analysis to understand the behavior of the eigenvalues of heavy-tailed Wigner matrices. We thus introduce several results on the eigenvalues of $M$ in the next section, before we analyze the integral in the right side of (2.4).
3 Preliminaries
3.1 Properties of heavy-tailed Wigner matrices
In this section, we collect several important results on heavy-tailed random variables and heavy-tailed Wigner matrices. For the sake of convenience, throughout this section we assume that $M$ is an heavy-tailed Wigner matrix with exponent $\alpha$ as in Definition 2.4 and each entry $M_{ij}$ satisfies $\mathbb{P}(|M_{ij}|>u)=G(u)=L(u)u^{-\alpha}$, where $L$ is a slowly varying function as in Definition 2.2. We let $\lambda_{1}\geq\dots\geq\lambda_{N}$ be the eigenvalues of $M$. We first define the high probability events.
Definition 3.1 (High probability event).
We say that an event $F_{N}$ holds with high probability if there exist (small) $\delta>0$ and (large) $N_{0}$ such that $\mathbb{P}(F_{N})<N^{-\delta}$ holds for any $N>N_{0}$.
The distribution of the largest eigenvalue $\lambda_{1}$ of $M$ is given as follows:
Proposition 3.2.
The limiting distribution of the rescaled largest eigenvalue $\lambda_{1}$ is given by
$\lim_{N\to\infty}\mathbb{P}\left(\frac{\lambda_{1}}{b_{N}}\leq x\right)=\exp(-x^{-\alpha}),$
where the normalization factor $b_{N}$ is as given in (2.1).
For the proof of Proposition 3.2, see Theorem 1 in [24]. Next, we state the results that ensure good events for $M$ hold with high probability.
Lemma 3.3.
With high probability, the following hold:
  • For any $1\leq i\leq N$, $|M_{ii}|\leq b_{N}^{11/20}$
  • For any $1\leq i\leq j\leq N$, we have $|M_{ij}|\leq b_{N}^{99/100}$ or $|M_{ii}|+|M_{jj}|\leq b_{N}^{1/10}$
  • For any constant $\delta>0$, there is no row of $M$ that has at least two entries whose absolute values are larger than $b_{N}^{1/2+\delta}$.
See [24] for the proof.
Lemma 3.4.
For every $\epsilon>0$, with high probability, that the following holds for any $i=1,2,\dots,N$: Suppose that $|M_{ii^{\prime}}|=\max_{j}|M_{ij}|$. Then, $|M_{ii^{\prime}}|<b_{N}^{1/2+\epsilon}$ or $|\sum_{j:j\neq i^{\prime}}M_{ij}|<b_{N}^{1/2+\epsilon}$.
For the proof, see Appendix A. As a simple corollary to Lemma 3.4, we can also obtain the following estimates on the extreme eigenvalues of $M$.
Lemma 3.5.
With high probability, we have
$\lambda_{1}=(1+O(b_{N}^{-1/8}))\max\{|M_{ij}|\},\quad|\lambda_{N}|=(1+O(b_{N}^{-1/8}))\max\{|M_{ij}|\}$
See Appendix A for the proof. We also can show that $\lambda_{1}$ is away from $\lambda_{2}$ and 0 with high probability. This will be used to show approximate the value of steepest descent method.
Lemma 3.6.
For every $\epsilon>0$, these three conditions hold with high probability.
  • $\lambda_{1}>b_{N}N^{-\epsilon}$
  • $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$
  • If $\lambda_{1}<Cb_{N}$, then $\lambda_{1}<Cb_{N}-b_{N}N^{-\epsilon}$
See Appendix A for the proof, where the assumption for slowly varying functions in Definition 2.2 plays a role. The next result is about the sum of i.i.d. heavy-tailed random variables.
Proposition 3.7.
Suppose $X_{1},X_{2},\cdots$ are i.i.d heavy-tailed random variables with exponent $\alpha$ defined in Definition 2.2 such that
$\lim_{x\to\infty}\mathbb{P}(X_{1}>x)/\mathbb{P}(|X_{1}|>x)=\theta\in[0,1]$
Let $S_{N}=X_{1}+\cdots+X_{N}$, $a_{N}=\inf\{x:\mathbb{P}(|X_{1}|>x)\leq N^{-1}\}$, and $c_{N}=N\mathbb{E}[X_{1}\mathbbm{1}_{(|X_{1}|\leq a_{N})}]$. Then, there exists a random variable $Y$ whose distribution is non-degenerate such that $(S_{N}-c_{N})/a_{N}\Rightarrow Y$.
For the proof, see Theorem 3.8.2 in [11].
Remark 3.8.
In Proposition 3.7, we can actually prove that both $c_{N}/a_{N}$ and $S_{N}/a_{N}$ converge. To see this, for $\alpha<1$, we consider
$\mathbb{E}[|X_{1}|\mathbbm{1}_{|X_{1}|\leq a_{N}}]=\int_{0}^{a_{N}}xg(x)\,\mathrm{d}x=\left.-xG(x)\right|_{0}^{a_{N}}+\int_{0}^{a_{N}}G(x)\,\mathrm{d}x$
where $g$ is the probability density function for $|X_{1}|$ and $G(x)=\mathbb{P}(|X_{1}|>x)$. The last term in the right side of the equation above is then
$\int_{a}^{a_{N}}G(x)\,\mathrm{d}x=\int x^{-\alpha}L(x)\,\mathrm{d}x=\frac{1}{1-\alpha}a_{N}^{-\alpha+1}L(a_{N})(1+o(1))=\frac{1}{1-\alpha}\frac{a_{N}}{N}(1+o(1))$
by Karamata’s theorem. (See, e.g., Thm 8.9.2 of [13]). Thus
$N\mathbb{E}[|X_{1}|\mathbbm{1}_{|X_{1}|\leq a_{N}}]/a_{N}\to\frac{1}{1-\alpha}-1,$
which implies that $c_{N}/a_{N}$ converges in probability to a constant depending only on the probability density function. Together with Proposition 3.7, this also shows that $S_{N}/a_{N}$ converges.
We can also prove the following bound for $a_{n}$ in Proposition 3.7, which will be used to prove an asymptotic behavior of the eigenvalue statistics.
Lemma 3.9.
Suppose that $X$ is a heavy-tailed random variables with exponent $\alpha$. Let $a_{N}=\inf\{x:\mathbb{P}(|X|>x)\leq N^{-1}\}$. Then, for any $\epsilon>0$, the following holds for any sufficiently large $N$:
$N^{1/\alpha-\epsilon}\leq a_{N}\leq N^{1/\alpha+\epsilon}.$
In particular, for any $\epsilon>0$,
$\lim_{N\to\infty}a_{N}/N^{1/\alpha+\epsilon}=0$
See Appendix A for the proof.
In the following lemmas, we prove the asymptotic behavior of $\sum\lambda_{i}^{2}$. We begin by considering the following estimate for the moments of $X$.
Lemma 3.10.
Suppose that the assumptions in Lemma 3.9 holds. Assume further that $X$ is non-negative. Then, for any $0<\delta<\alpha$, we have
$\mathbb{E}(X^{\delta})=O(1)$
See Appendix A for the proof. Lemma below is just corollary of this lemma.
Lemma 3.11.
For non-negative heavy-tailed random variables $X_{1},...,X_{N}$ with exponent $\alpha$,
$\mathbb{E}((\sum_{i=1}^{N}X_{i})^{\delta})=O(N)$
for $\delta<\alpha,1$.
Proof.
Jensen’s inequality implies $\mathbb{E}((\sum_{i=1}^{N}X_{i})^{\delta})\leq\sum_{i=1}^{N}\mathbb{E}(X_{i}^{\delta})=O(N).$
We next show that $\sum_{i}\lambda_{i}^{2}<N^{4/\alpha+\epsilon}$ holds with high probability.
Lemma 3.12.
For eigenvalues $\lambda_{1},\cdots,\lambda_{N}$ of $M$,
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>N^{4/\alpha+\epsilon})=O(N^{-\alpha\epsilon/4}).$
In particular,
$\sum_{i}\lambda_{i}^{2}<N^{4/\alpha+\epsilon}$
with high probability.
See Appendix A for the proof. The next lemma provides the information about the asymptotic for the number of eigenvalues that is comparable with the largest eigenvalue.
Lemma 3.13.
For the eigenvalues $\lambda_{1},\cdots,\lambda_{N}$ of $M$,
#$\#\{|\lambda_{i}|>b_{N}N^{-\epsilon}\}=O(N^{3\epsilon})$
with high probability.
See Appendix A for the proof. Furthermore, we can prove that $\sum\lambda_{i}^{2}/b_{N}^{2}$ converges in distribution to a random variable, which will be used to show that our free energy also converges in distribution to a random variable.
Lemma 3.14.
The eigenvlaues $\lambda_{1},..,\lambda_{N}$ of M satisfy
$\sum_{i}\lambda_{i}^{2}/b_{N}^{2}\Rightarrow X$
for some non-degenerate random variable $X$.
See Appendix A for the proof.
Furthermore, we also need the limit of $\sum_{i}\lambda_{i}/b_{N}$ and it converges to 0 as below lemma.
Lemma 3.15.
The eigenvlaues $\lambda_{1},..,\lambda_{N}$ of $M$ satisfies
$\lim_{N\to\infty}\sum_{i=1}^{N}\lambda_{i}/b_{N}=0.$
See Appendix A for the proof.
3.2 Integral representation formula for the partition function
Lemma 3.16.
Let $Z_{N}$ be the partition function defined in Definition 2.5. Then,
$Z_{N}=C_{N}\frac{1}{i}\int_{\gamma-i\infty}^{\gamma+i\infty}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z$ (3.1)
for any $\gamma>\lambda_{1}$, where
$G(z)=2\beta z-\frac{b_{N}}{N}\sum_{i}\log(z-\lambda_{i}),\quad C_{N}=\frac{\Gamma(N/2)b_{N}^{N/2-1}}{2\pi(N\beta)^{N/2-1}}.$ (3.2)
The proof of Lemma 3.16 is standard (see, e.g., [5]). For the detailed proof, see Appendix A.
Since $G^{\prime}$ is an incresing function on the interval $(\lambda_{1},\infty)$ with
$\lim_{z\to\lambda_{1}^{+}}G^{\prime}(z)=-\infty,\quad\lim_{z\to\infty}G^{\prime}(z)=2\beta>0,$
it is immediate to see that there exists a unique solution to the equation $G^{\prime}(z)=0$, which with abuse of notation we call $\gamma$ in this section. In the following lemma, we prove a result on the relation between the critical point $\gamma$ and $\lambda_{1}$.
Lemma 3.17.
Set $\gamma\in(\lambda_{1},\infty)$ be the (unique) solution to the equation $G^{\prime}(z)=0$. Then,
  • If $\lambda_{1}<\frac{b_{N}}{2\beta}$, there exists a random variable $X_{N}$ such that
    $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N})$
    and $X_{N}$ converges in distribution to a non-degenerate random variable $X$ as $N\to\infty$.
  • If $\lambda_{1}>\frac{b_{N}}{2\beta}$, for any $80<\epsilon<1/$ 8,
    $\gamma=\lambda_{1}+\frac{1}{2\beta-b_{N}/\lambda_{1}}\frac{b_{N}}{N}+O(\frac{b_{N}}{N^{1+\epsilon}})$
    with high probability.
For the proof, we can simply approximate $\gamma$ appropriately with some steps. See Appendix A for the proof.
4 High temperature regime
In this section, we consider the regime $\lambda_{1}<b_{N}/2\beta$. Recall that we let $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N}+O(N^{-1-\epsilon}))$ in Lemma 3.17.
Proposition 4.1.
For event $F_{1}=\left\{\lambda_{1}\leq\frac{b_{N}}{2\beta}\right\}$,
$(\log Z_{N})|F_{1}\Rightarrow T$
for some random variable $T$. Moreover, $\mathbb{P}(F_{1})\to\exp(-(2\beta)^{\alpha})$ as $N\to\infty$.
Due to Lemma 3.5, we have $\lambda_{1}=\max\{|M_{ij}|\}(1+O(b_{N}^{-1/8})),\lambda_{N}=-\max\{|M_{ij}|\}(1+O(b_{N}^{-1/8}))$ holds with high probability. Hence, for $F_{1}^{\prime}=\{|M_{ij}|\leq\frac{b_{N}}{2\beta}\ \textit{for all }i,j\}$, $\mathbb{P}(F_{1}^{\prime}\Delta F_{1})\leq N^{-\delta}$ for some $\delta>0$. Also, for $F_{1}^{\prime\prime}=\{|M_{ij}|\leq\frac{b_{N}}{2\beta}(1-N^{-\epsilon})\textit{for all }i,j\}$ for small $\epsilon>0$, then $\mathbb{P}(F_{1}^{\prime\prime}\Delta F_{1})\leq N^{-\delta}$ for some $\delta>0.$ Therefore, we condition the matrix and all the statistics to event $F_{1}^{\prime\prime}$ in this section. For the convenience, we express variables same as before. Similar to the low temperature case, we first prove the following lemma for the derivatives of the function $G.$
Lemma 4.2.
With high probability, the k-th derivative of G satisfies
$G^{(k)}(a)=O(b_{N}^{-k+1})$
where $|\gamma-a|=O(b_{N}N^{-\delta})$ for some $\delta>0$.
Proof.
Lemma 3.13 implies that
#$\#\{|\lambda_{i}|>b_{N}N^{-\epsilon}\}=O(N^{3\epsilon})$
with high probability and Lemma 3.6 implies $\lambda_{1}<\frac{b_{N}}{2\beta}(1-N^{-\epsilon})$ holds for every $\epsilon>0$ from $\lambda_{1}<\frac{b_{N}}{2\beta}$. Since $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N})$ from Lemma 3.17, we conclude that
$\begin{split}G^{(k)}(a)&=\frac{b_{N}}{N}(k-1)!\sum_{i=1}^{N}\frac{1}{(a-\lambda_{i})^{k}}=\frac{b_{N}(k-1)!}{N}\left(O(N^{3\epsilon})\frac{N^{k\epsilon}}{b_{N}^{k}}+(N-O(N^{3\epsilon}))O({b_{N}^{-k}})\right)\\ &=O(b_{N}^{-k+1})\end{split}$
with high probability. ∎
Proof of Proposition 4.1.
Recall that the free energy is
$\log Z_{N}=\log C_{N}+\frac{N}{2b_{N}}G(\gamma)+\log\int_{-\infty}^{+\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t.$
To handle the last term, we consider the following representation with a sufficiently small $\epsilon>0,$
$\int_{-\infty}^{+\infty}$ $\exp\left({\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\right)dt=\sqrt{\frac{b_{N}}{N}}\int_{-\infty}^{+\infty}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t$
$=\sqrt{\frac{b_{N}}{N}}\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t$ (4.1)
$~{}~{}+\sqrt{\frac{b_{N}}{N}}\int_{|t|>\sqrt{b_{N}}N^{\epsilon}}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t.$ (4.2)
From Lemma 4.2 and Taylor approximation,
(4.1)$&=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+O(\frac{N}{2b_{N}}\frac{1}{3!}\sup_{|a-\gamma|<b_{N}N^{\epsilon-1/2}}|G^{(3)}(a)|(\sqrt{b_{N}}t/\sqrt{N})^{3}}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+\frac{N}{2b_{N}}O(b_{N}^{-2}(b_{N}N^{\epsilon-1/2})^{3})}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+O(N^{-\epsilon})}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}e^{(-\frac{1}{4}G^{\prime\prime}(\gamma))t^{2}}\mathrm{d}t(1+O(N^{-\epsilon})).\end{split}$ (4.3)
Since $G^{\prime\prime}(\gamma)=O(b_{N}^{-1})$ and
$\int_{x\geq a}e^{-\lambda x^{2}}\mathrm{d}x=\int_{0}^{\infty}e^{-\lambda(t+a)^{2}}\mathrm{d}t=e^{-\lambda a^{2}}\int_{0}^{\infty}e^{-\lambda(2ta+t^{2})}\mathrm{d}t\leq e^{-\lambda a^{2}}\int_{0}^{\infty}e^{-\lambda t^{2}}\mathrm{d}t=\sqrt{\frac{\pi}{2\lambda}}e^{-\lambda a^{2}},$
we obtain
$\int_{-\sqrt{b_{N}}N^{\epsilon}}^{\sqrt{b_{N}}N^{\epsilon}}e^{-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}}\mathrm{d}t=\int_{-\infty}^{+\infty}e^{-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}}\mathrm{d}t+O(e^{-N^{\epsilon}})=2\sqrt{\frac{\pi}{G^{\prime\prime}(\gamma)}}+O(e^{-N^{\epsilon}}).$
For the bound of (4.2), we have
|(4.2)$|&=\left|\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\operatorname{\mathrm{Re}}\exp\left({\frac{N}{2b_{N}}(2\beta ti\sqrt{\frac{b_{N}}{N}}-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(1-\frac{\sqrt{\frac{b_{N}}{N}}ti}{\gamma-\lambda_{i}}))}\right)\mathrm{d}t\right|\\ &\leq\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\left|\exp\left({\frac{N}{2b_{N}}(2\beta ti\sqrt{\frac{b_{N}}{N}}-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(1-\frac{\sqrt{\frac{b_{N}}{N}}ti}{\gamma-\lambda_{i}}))}\right)\right|\mathrm{d}t\\ &=\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\exp\left({-\frac{1}{4}\sum_{i=1}^{N}\log(1+\frac{{\frac{b_{N}}{N}}t^{2}}{(\gamma-\lambda_{i})^{2}})}\right)\mathrm{d}t.\end{split}$ (4.4)
Note that there are at least $N/2$ eigenvalues that are less than equal to $b_{N}$ with high probability, since $\sum\lambda_{i}^{2}<b_{N}^{2}N^{\epsilon}$ with high probability. Using this fact, we further have
[…]
$\leq\int_{\sqrt{b_{N}}N^{\epsilon}}^{b_{N}N}\exp\left({-\frac{b_{N}N^{2\epsilon}}{8N}}\right)\mathrm{d}t+\int_{t>b_{N}N}2\left(\frac{t^{2}}{b_{N}N}\right)^{-N/8}\mathrm{d}t$
$=O(b_{N}Ne^{-N^{2\epsilon}})+\left.\frac{1}{1-N/4}t^{-N/4+1}(c^{2}b_{N}N)^{N/8}\right|_{t=b_{N}N}^{t=\infty}$
$=O(b_{N}Ne^{-N^{2\epsilon}})+O(N^{-N/8}).$
Hence,
$\log\int_{-\infty}^{\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t=\frac{1}{2}\log(b_{N}/N)+\frac{1}{2}\log(\frac{4\pi}{G^{\prime\prime}(\gamma)})+O(N^{-\epsilon}).$
On the other hand,
$G^{\prime\prime}(\gamma)=\frac{b_{N}}{N}\sum_{i=1}^{N}\frac{1}{(\gamma-\lambda_{i})^{2}}$
and
$\sum_{i=1}^{N}\frac{1}{(\gamma-\lambda_{i})^{2}}=O(N^{3\epsilon})\frac{1}{N^{-2\epsilon}b_{N}^{2}}+(N-O(N^{3\epsilon}))\frac{1+O(N^{-\epsilon})}{\gamma^{2}}=\frac{N}{(b_{N}/2\beta)^{2}}(1+O(N^{-\epsilon})).$
Hence,
$\log G^{\prime\prime}(\gamma)=\log\frac{(2\beta)^{2}}{b_{N}}+O(N^{-\epsilon}).$
Putting these results together, we have
$\log Z_{N}=\log C_{N}+\frac{N}{2b_{N}}G(\gamma)+\log\int_{-\infty}^{+\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t$
$=\log C_{N}+\frac{N\beta}{b_{N}}\gamma-\frac{1}{2}\sum_{i=1}^{N}\log(\gamma-\lambda_{i})+\frac{1}{2}\log(b_{N}/N)+\frac{1}{2}\log(\frac{4\pi}{G^{\prime\prime}(\gamma)})+O(N^{-\epsilon})$
$=\log C_{N}+\frac{N}{2}-\frac{N}{2}\log(\frac{b_{N}}{2\beta})+\frac{1}{2}\log(b_{N}/N)+\frac{1}{2}\log(\frac{4\pi}{(2\beta)^{2}/b_{N}})$ (4.5)
$~{}~{}~{}+\frac{N\beta}{b_{N}}(\gamma-\frac{b_{N}}{2\beta})-\frac{N}{2}\log\frac{\gamma}{b_{N}/2\beta}+O(N^{-\epsilon})-\frac{1}{2}\sum_{i=1}^{N}\log(1-\frac{\lambda_{i}}{\gamma}).$ (4.6)
For the first term in the right side,
(4.5)$=\log\frac{b_{N}^{N/2-1}\Gamma(N/2)}{2\pi(N\beta)^{N/2-1}}+\frac{N}{2}-\frac{N}{2}\log(\frac{b_{N}}{2\beta})+\frac{1}{2}\log(\frac{4\pi b_{N}^{2}}{(2\beta)^{2}N})$ )
$~{}~{}=\log\frac{\Gamma(N/2)}{2\pi N^{N/2-1}}+\frac{N}{2}+\frac{N}{2}\log 2+\frac{1}{2}\log(\frac{\pi}{N})$
$~{}~{}=\log\frac{\sqrt{\frac{4\pi}{N}}(\frac{N}{2e})^{N/2}}{2\pi N^{N/2-1}}+\frac{N}{2}\log(2e)+\frac{1}{2}\log(\pi/N)+O(N^{-1})=O(N^{-1}),$
since Stirling’s formula implies
$\Gamma(N/2)=\sqrt{\frac{4\pi}{N}}\left(\frac{N}{2e}\right)^{N/2}(1+O(1/N)),$ $\log\Gamma(N/2)=\log\sqrt{\frac{4\pi}{N}}\left(\frac{N}{2e}\right)^{N/2}+O(1/N).$
We next consider the term (4.6). The first two terms of (4.6) can be simplified to
$\frac{N\beta}{b_{N}}(\gamma-\frac{b_{N}}{2\beta})-\frac{N}{2}\log(1+\frac{\gamma-b_{N}/2\beta}{b_{N}/2\beta})=\frac{X_{N}}{2}-\frac{N}{2}\log(1+\frac{X_{N}}{N}),$
and it converges in probability to 00, which can be checked by an elementary calculation
$\lim_{N\to\infty}\left(\frac{x}{2}-\frac{N}{2}\log(1+\frac{x}{N})=0\right)$
for all $x>0$. Furthermore, in Lemma 4.3, we prove that $-\frac{1}{2}\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution to some random variable $T$. In conclusion, if we define $F_{N}:=\log Z_{N}$, then it converges in distribution to a random variable $T$. ∎
In Appendix B, we prove several the statistical properties of $T$.
We conclude this section by proving that $T_{N}=-\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution.
Lemma 4.3.
The log statistics $T_{N}=-\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution to some random variable.
Proof.
From Lemma 3.5,
$|\lambda_{1}|=\max_{i,j}\{|M_{ij}|\}(1+O(b_{N}^{-1/8})),\ \ \ |\lambda_{n}|=\max_{i,j}\{|M_{ij}|\}(1+O(b_{N}^{-1/8}))$
with high probability, and hence $|\lambda_{n}|,\lambda_{1}<\gamma$ with high probability. Thus,
$-\log(1-\lambda_{i}/\gamma)=\sum_{k=0}^{\infty}\frac{1}{k}(\frac{\lambda_{i}}{\gamma})^{k}.$
Let
$S_{k}=\sum_{i=1}^{N}(\frac{\lambda_{i}}{\gamma})^{k},$
then Lemma 3.15 implies
$S_{1}=\sum_{i=1}^{N}\lambda_{i}/\gamma=Tr(M_{N})/\gamma\to 0.$
Moreover, from Lemma 3.14
$S_{2}=\sum_{i=1}^{N}(\frac{\lambda_{i}}{\gamma})^{2}=\frac{\sum\lambda_{i}^{2}}{b_{N}^{2}}\frac{b_{N}^{2}}{\gamma^{2}}\Rightarrow X$
for some non-degenerate random variable $X$. Note that, for $k\geq 3$,
$|\sum_{i=1}^{N}(\frac{\lambda_{i}}{\gamma})^{k}|\leq|\frac{\Gamma}{\gamma}|^{k-2}S_{2}/\gamma^{2}$
where we let $\Gamma=\max\{|\lambda_{n}|,\lambda_{1}\}$. Thus,
$|T_{N}|\leq\sum_{k=3}^{\infty}\frac{1}{k}|\frac{\Gamma}{\gamma}|^{k-2}S_{2}\leq-(\frac{\gamma}{\Gamma})^{2}{S_{2}}\log(1-\Gamma/\gamma).$
Since $S_{2}$ and ${\Gamma}/{\gamma}$ converge in distribution to some random variables, this proves $T_{N}$ converges to a non-degenerate random variable. ∎
5 Low temperature regime
In this section, we assume the event $F_{2}=\left\{\lambda_{1}>\frac{b_{N}}{2\beta}\right\}$ holds. Our goal in this section is to prove the following proposition:
Proposition 5.1.
Suppose that assumptions in Theorem 2.6 hold. Then, given the event $F_{2}=\left\{\lambda_{1}>\frac{b_{N}}{2\beta}\right\}$,
$(\frac{1}{N}\log Z_{N})|F_{2}\Rightarrow\beta X-\frac{1}{2}\log(2e\beta X)$
where the random variable $X$ satisfies $X>\frac{1}{2\beta}$ and its distribution is given by
$\mathbb{P}(X>u)=\exp((2\beta)^{\alpha}-u^{-\alpha}),.$
Moreover, $\lim_{N\to\infty}\mathbb{P}(F_{2})=1-\exp(-(2\beta)^{\alpha})$.
We begin by proving the asymptotic behavior of the Taylor coefficients of the function $G$ defined in (2.5).
Lemma 5.2.
With high probability, the Taylor coefficient $G$ satisfies
$G^{(k)}(\gamma)=\Theta(N^{k-1}b_{N}^{-(k-1)}).$
Proof.
Recall that we proved $(\gamma-\lambda_{1})=\Theta(b_{N}/N)$ in Lemma 3.17. Then,
$G^{(k)}(\gamma)=\frac{b_{N}}{N}\sum_{i}\frac{(k-1)!}{(\gamma-\lambda_{i})^{k}}\geq\frac{b_{N}}{N}\frac{1}{(\gamma-\lambda_{1})^{k}}=\Theta(N^{k-1}b_{N}^{-k+1}).$
Since $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$ with high probability due to Lemma 3.6,
$\begin{split}G^{(k)}(\gamma)&=\frac{b_{N}}{N}\sum_{i}(k-1)!\frac{1}{(\gamma-\lambda_{i})^{k}}\\ &=O\left(\frac{b_{N}}{N}\frac{1}{(\gamma-\lambda_{1})^{k}}\right)+O\left(N\frac{b_{N}}{N}b_{N}^{-k}N^{k\epsilon}\right)=O(N^{k-1}b_{N}^{-k+1}).\end{split}$
We thus conclude that
$G^{(k)}(\gamma)=\Theta(N^{k-1}b_{N}^{-k+1}).$
We now consider the logarithm of the partition function,
$\log Z_{N}=\log C_{N}+\frac{N}{2b_{N}}G(\gamma)+\log\int_{-\infty}^{+\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t.$
We claim that the last term in the equation above satisfies
$\lim_{N\to\infty}\frac{1}{N}\log\int_{-\infty}^{+\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t=0.$
To prove this, we will show
$C_{1}N^{c_{1}}\leq\int_{-\infty}^{+\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t\leq C_{2}N^{c_{2}}$
for some $C_{1},C_{2}>0$ and $c_{1},c_{2}\in\mathbb{R}$. For the upper bound, we have
$\left|\int_{-\infty}^{+\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t\right|=\left|\int_{0}^{+\infty}2\operatorname{\mathrm{Re}}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\mathrm{d}t\right|$
$\leq\int_{0}^{+\infty}2\left|e^{\frac{N}{2b_{N}}(2\beta ti)-\frac{1}{2}\sum\log(1+\frac{ti}{\gamma-\lambda_{i}})}\right|\mathrm{d}t=\int_{0}^{+\infty}2e^{-\frac{1}{4}\sum\log(1+\frac{t^{2}}{(\gamma-\lambda_{i})^{2}})}\mathrm{d}t$
$\leq\int_{0}^{b_{N}^{4}}1\mathrm{d}t+\int_{b_{N}^{4}}^{\infty}\left(\frac{t^{2}}{\lambda_{1}^{2}}\right)^{-N/8}\mathrm{d}t=O(b_{N}^{4}),$
where we used the Taylor expansion of $G(\gamma+t)$ with Lemma 5.2 as
$G(\gamma+t)=\sum\frac{1}{k!}G^{(k)}(\gamma)t^{k}=\sum O(b_{N}^{-(k-1)}N^{(k-1)})t^{k}=G(\gamma)+\frac{b_{N}}{N}\sum_{k\geq 2}O((\frac{N}{b_{N}}t)^{k}),$
which converges for $|t|<\frac{b_{N}}{N}$.
The remaining part is to show its lower bound. To show its lower bound, we will use contour integral. Consider the curve $\Gamma$ which satisfies $\operatorname{\mathrm{Im}}(G(t))=0$ on that curve and starting from $\gamma$ and end with $\infty$- ∞. By the Cauchy-Riemann equation $\operatorname{\mathrm{Re}}(G(z))$ is monotone along the curve. Since $G(z)$ decreases near $\gamma$ as imaginary part increases from its Taylor expansion, $G(z)=\operatorname{\mathrm{Re}}(G(z))$ decays along the curve.
We consider the path $\Gamma^{+}=\Gamma\cap\mathbb{C}^{+}$ and $\Gamma^{-}=\Gamma\cap\mathbb{C}^{-}$ which are starting from $\gamma$ and goes to $\infty$- ∞. Let $C_{R}=\{|z|=R\}\cap C^{+}\cap\operatorname{\mathrm{Re}}(z)<\gamma$, then $\operatorname{\mathrm{Re}}(G(z))\leq 2\beta\gamma-b_{N}\log(R/2)$
$|\int_{C_{R}}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z|\leq\frac{T_{N}}{R^{N/2}-1}$
for some constant $T_{N}$ which only depends on $N$.
This implies
$\int_{\gamma-i\infty}^{\gamma+i\infty}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z=\int_{\Gamma}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z=\int_{\Gamma^{+}}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z+\int_{\Gamma^{-}}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z=2\int_{\Gamma^{+}}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z,$
since $\overline{G(\bar{z})}=G(z)$.
Let
$K=-i\int_{\gamma-i\infty}^{\gamma+i\infty}\exp(\frac{N}{2b_{N}}(G(z)-G(\gamma)))\mathrm{d}z=-2i\int_{\Gamma^{+}}\exp(\frac{N}{2b_{N}}[G(z)-G(\gamma)])\mathrm{d}z,$
then
$K=-2i\int_{\Gamma^{+}}\exp(\frac{N}{2b_{N}}[G(z)-G(\gamma)])\mathrm{d}x+2\int_{\Gamma^{+}}\exp(\frac{N}{2b_{N}}[G(z)-G(\gamma)])\mathrm{d}y.$
To obtain the lower bound of K, we will prove the lower bound for the latter real part. Its Taylor series holds for $|t|<\frac{b_{N}}{N}$, and
$G(\gamma+t)=G(\gamma)+\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)t^{k}.$
On the curve $\Gamma^{+}$
$0=\operatorname{\mathrm{Im}}G(\gamma+x+iy)=xyG^{\prime\prime}(\gamma)-\frac{1}{6}G^{{}^{\prime\prime\prime}}(\gamma)y^{3}+\frac{1}{2}G^{{}^{\prime\prime\prime}}(\gamma)x^{2}y+\sum_{k\geq 4}\frac{1}{k!}G^{(k)}(\gamma)\operatorname{\mathrm{Im}}(x+iy)^{k}$
and
$|\operatorname{\mathrm{Im}}(x+iy)^{k}|\leq k|y||x+iy|^{k-1}.$
This implies below for $|t|\leq\frac{1}{N^{2}}$;
$\Omega(x,y)$ $=\frac{1}{G^{{}^{\prime\prime}}(\gamma)y}\sum_{k\geq 4}\frac{1}{k!}G^{(k)}(\gamma)\operatorname{\mathrm{Im}}(x+iy)^{k}\leq\sum_{k\geq 4}\frac{1}{(k-1)!}\frac{G^{(k)}(\gamma)}{G^{{}^{\prime\prime}}(\gamma)}|x+iy|^{k-1}$
$\leq O\left(\frac{N^{2}}{b_{N}^{2}}\right)(x^{2}+y^{2})|t|\leq O((\frac{N}{b_{N}})^{2}\frac{1}{N^{2}})(x^{2}+y^{2}).$
Set $\alpha=-\frac{G^{\prime\prime\prime}(\gamma)}{G^{\prime\prime}(\gamma)}>0$ and $\alpha=O(\frac{N}{b_{N}})$, then we have
00 0 $=x+\frac{1}{6}\alpha y^{2}-\frac{1}{2}\alpha x^{2}+O((\frac{N}{b_{N}})^{2}\frac{1}{N^{2}})(x^{2}+y^{2})$
$=x+\frac{1}{6}\alpha y^{2}(1+o(1))-\frac{1}{2}\alpha x^{2}(1+o(1))=x(1+o(1))+\frac{1}{6}\alpha y^{2}(1+o(1)).$
Therefore,
$x=-\frac{1}{6}\alpha y^{2}(1+o(1))\text{ for }|x+iy|\leq\frac{1}{N^{2}}$
and $\operatorname{\mathrm{Im}}y$ increases on $|t|<\frac{1}{N^{2}}$.
Since
$t^{2}=x^{2}+y^{2}=\{O(\frac{N}{b_{N}})y^{2}\}^{2}+y^{2}=O(\frac{N^{2}}{b_{N}^{2}})y^{4}+y^{2},$
we have
$y=\Theta(t),x=\Theta\left(\frac{N}{b_{N}}t^{2}\right)\text{ for }t<\frac{1}{N^{2}}.$
Now we will prove its lower bound using Lemma 5.3: for any real valued function f on $\Gamma^{+}$ which decreases along the curve as $z$ moves from $\gamma$ to $\infty$- ∞, $\int_{\Gamma^{+}}e^{f(z)}\mathrm{d}y\geq 0$. Hence, we construct below decreasing function and we show our integration is greater than their difference.
Let
$\Gamma_{1}=\{|z|\leq N^{-3}\}\cap\Gamma^{+},\Gamma_{2}=\{N^{-3}\leq|z|\leq N^{-2}\}\cap\Gamma^{+},\Gamma_{3}=\{|z|\geq N^{-2}\}\cap\Gamma^{+}$
$z_{3}=\{|z|=N^{-3}\}\cap\Gamma^{+},z_{2}=\{|z|=N^{-2}\}\cap\Gamma^{+}.$
Then we have
$\int_{\Gamma_{3}}\mathrm{d}y\geq\operatorname{\mathrm{Im}}z_{3}-\operatorname{\mathrm{Im}}\gamma\geq C_{3}N^{-3}.$
Let
$g(z)=\begin{cases}\frac{N}{2b_{N}}[G(z)-G(\gamma)],&|z|>N^{-2}\\ \frac{N}{2b_{N}}[G(z_{2})-G(\gamma)],&|z|<N^{-2}.\end{cases}$
Then $g(z)$ decreases along the curve $\Gamma^{+}$
$\int_{\Gamma^{+}}\exp(\frac{N}{2b_{N}}[G(z)-G(\gamma)])\mathrm{d}y\geq\int_{\Gamma_{1}}\exp(\frac{N}{2b_{N}}[G(z_{3})-G(\gamma)])\mathrm{d}y+\int_{\Gamma_{2}\cup\Gamma_{3}}\exp(g(z)))\mathrm{d}y$
$\geq\int_{\Gamma_{1}}\exp(\frac{N}{2b_{N}}[G(z_{3})-G(\gamma)])-\exp(\frac{N}{2b_{N}}[G(z_{2})-G(\gamma)])\mathrm{d}y+\int_{\Gamma^{+}}\exp({g(z)})\mathrm{d}y$
$\geq\int_{\Gamma_{1}}\exp(\frac{N}{2b_{N}}[G(z_{3})-G(\gamma)])-\exp(\frac{N}{2b_{N}}[G(z_{2})-G(\gamma)])\mathrm{d}y.$
Since
$|G(\gamma+t)-G(\gamma)|=|\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)t^{k}|=|\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)t^{k}|=\Theta(\frac{N}{b_{N}}t^{2})$
for $|t|<N^{-2}$ and $G(\gamma+t)<G(\gamma)$ along the curve,
$\int_{\Gamma_{1}}\exp(\frac{N}{2b_{N}}[G(z_{3})-G(\gamma)])-\exp(\frac{N}{2b_{N}}[G(z_{2})-G(\gamma)])\mathrm{d}y$
$\geq\int_{\Gamma_{1}}\exp(-C_{1}\frac{N}{2b_{N}}N^{-4})-\exp(-C_{2}\frac{N}{b_{N}}N^{-6})\mathrm{d}y\geq C_{3}N^{-3}\frac{N}{2b_{N}}N^{-4}.$
Therefore, this proves that $\frac{1}{N}\log\int_{-\infty}^{\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\to 0$.
Lemma 5.3.
Let $f$ be a real-valued function defined on $\Gamma^{+}$ and $f(z)$ is decreasing along the curve $\Gamma^{+}$ as z moves from $\gamma$ to $\infty$- ∞
$\int_{\Gamma^{+}}e^{f(z)}\mathrm{d}y\geq 0.$
Proof of Lemma 5.3.
Consider the function $s:[0,1]\to\mathbb{C}$ such that $s(0)=\gamma$, $s(1)=-\infty$ and which follows the contour of $\Gamma^{+}$, then we have
$\int_{\Gamma^{+}}e^{f(z)}\mathrm{d}y=\int_{0}^{1}e^{f(s(t))}\operatorname{\mathrm{Im}}s^{\prime}(t)\mathrm{d}t.$
As $t$ moves from 0 to 1, we can divide the interval $[0,1]$ into $0=s_{0}<s_{1}<...,<s_{k}=1$ such that $\operatorname{\mathrm{Im}}s(t)$ increases on $[s_{2m},s_{2m+1}]$ and decreases on $[s_{2m+1},s_{2m+2}]$.  Let $\Delta_{k}=s(s_{k+1})-s(s_{k})$,then we have
$\int_{s_{2m}}^{s_{2m+1}}e^{f(s(t))}\operatorname{\mathrm{Im}}s^{\prime}(t)\mathrm{d}t\geq e^{s_{2m+1}}(s(s_{2m+1})-s(s_{2m})),$
$\int_{s_{2m-1}}^{s_{2m}}e^{f(s(t))}\operatorname{\mathrm{Im}}s^{\prime}(t)\mathrm{d}t\geq-e^{s_{2m-1}}(s(s_{2m})-s(s_{2m-1})).$
Therefore,
$\int_{\Gamma^{+}}e^{f(z)}\mathrm{d}y$ $=\int_{0}^{1}e^{f(s(t))}\operatorname{\mathrm{Im}}s^{\prime}(t)\mathrm{d}t\geq\sum e^{s_{2m+1}}(\Delta_{2m}-\Delta_{2m+1})$
$\geq e^{f(s_{1})}(\Delta_{0}-\Delta_{1})+\sum_{m\geq 1}e^{s_{2m+1}}(\Delta_{2m}-\Delta_{2m+1})$
$\geq e^{f(s_{1})}(\Delta_{0}-\Delta_{1})+e^{f(s_{3})}(\Delta_{2}-\Delta_{3})+\sum_{m\geq 2}e^{s_{2m+1}}(\Delta_{2m}-\Delta_{2m+1})$
$\geq e^{f(s_{3})}((\Delta_{0}-\Delta_{1})+(\Delta_{2}-\Delta_{3}))+\sum_{m\geq 2}e^{s_{2m+1}}(\Delta_{2m}-\Delta_{2m+1})$
$\geq e^{f(s_{2k+1})}(\sum_{m=0}^{k}\Delta_{2m}-\Delta_{2m+1})+\sum_{m\geq k+1}e^{s_{2m+1}}(\Delta_{2m}-\Delta_{2m+1})$
$\geq e^{f(s_{2t+1})}\sum\Delta_{2m}-\Delta_{2m+1}\geq 0.$
This comes from the fact that $\Gamma^{+}$ does not across the real line and it implies $\sum_{m=0}^{k}\Delta_{2m}-\Delta_{2m+1}\geq 0$ for arbitrary $k\geq 0$.
Now, let us get back to the proof of Proposition 5.1.
Proof of Proposition 5.1.
The logarithm of partition function
$\log Z_{N}=\log C_{N}+\frac{N\beta\gamma}{b_{N}}-\frac{1}{2}N\log b_{N}-\frac{1}{2}\sum_{i=1}^{N}\log(\gamma/b_{N}-\lambda_{i}/b_{N})+\log\int_{-\infty}^{\infty}e^{\frac{N}{2b_{N}}(G(\gamma+it)-G(\gamma)}\mathrm{d}t.$
Since we proved
$\lim_{N\to\infty}\frac{1}{N}\log\int_{-\infty}^{\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}=0,$
the limiting distribution of the free energy $F_{N}=\frac{1}{N}\log Z_{N}$ is equal to that of
$\frac{1}{N}\log C_{N}-\frac{1}{2}\log b_{N}+\frac{\beta\gamma}{b_{N}}-\frac{1}{2N}\sum_{i=1}^{N}\log(\gamma/b_{N}-\lambda_{i}/b_{N})$
$=\frac{1}{N}\left[\log\frac{\Gamma(N/2)b_{N}^{N/2-1}}{2\pi(N\beta)^{N/2-1}b_{N}^{N/2}}-\frac{1}{2}\sum_{i=1}^{N}\log(\gamma/b_{N}-\lambda_{i}/b_{N})\right]+\frac{\beta\gamma}{b_{N}}.$
Stirling’s approximation implies
$\Gamma(N/2)=\sqrt{\frac{4\pi}{N}}\left(\frac{N}{2e}\right)^{N/2}(1+O(1/N)),$
hence
$\lim\frac{1}{N}\log\frac{\Gamma(N/2)}{2\pi(N\beta)^{N/2-1}b_{N}}=-\frac{1}{2}\log(2e\beta).$
On the other hand,
$\sum\log(\gamma/b_{N}-\lambda_{i}/b_{N})$ $=\sum_{|\lambda_{i}|<b_{N}N^{-\epsilon}}\log(\gamma/b_{N}-\lambda_{i}/b_{N})+\sum_{|\lambda_{i}|>b_{N}N^{-\epsilon}}\log(\gamma/b_{N}-\lambda_{i}/b_{N})$
$=(N-O(N^{3\epsilon}))\log(\gamma/b_{N})(1+O(N^{-\epsilon}))+O(N^{3\epsilon})O(\log N).$
This implies that the limit of $\frac{1}{N}\sum_{i=1}^{N}\log(\gamma/b_{N}-\lambda_{i}/b_{N})$ is governed by the leading order term $\log(\lambda_{1}/b_{N}).$ Together with all above results, the limiting distribution of free energy
$F_{N}$ $=\frac{1}{N}\log\frac{\Gamma(N/2)}{2\pi(N\beta)^{N/2-1}b_{N}}-\frac{1}{2N}\sum_{i=1}^{N}\log(\gamma/b_{N}-\lambda_{i}/b_{N})+\frac{\beta\lambda_{1}}{b_{N}}$ (5.1)
$+\frac{1}{N}\log\int_{-\infty}^{\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}$
is equal to
$-1/2\log(2e\beta\lambda_{1}/b_{N})+\frac{\beta\lambda_{1}}{b_{N}}\Rightarrow-\frac{1}{2}\log\beta X+\beta X-1/2\log(2e),$ (5.2)
where $X$ is the limiting distribution of $\frac{\lambda_{1}}{b_{N}}.$ We conclude the proof of the desired statement. ∎
Appendix A Proofs of technical lemmas
Proof of Lemma 3.4.
We consider the cases $21<\alpha<$ 2 and $10<\alpha<$ 1 separately.
Case 1: $21<\alpha<$ 2.
We choose $\epsilon=\frac{1}{2M+1}$, where M is an integer satisfying $\epsilon<\frac{1}{4}-\frac{\alpha}{8}$. Then,
$\mathbb{E}(\#\{1\leq j\leq N:|M_{ij}|\geq b_{N}^{k\epsilon}\})=NG(b_{N}^{k\epsilon}).$
For binomial random variable $X\sim B(N,p)$, Chernoff’s inequality implies $\mathbb{P}(X\geq\mathbb{E}(X)+t)\leq\exp(-\frac{t^{2}}{2Np}+\frac{t^{3}}{6(Np)^{2}})$. This implies
(#$\mathbb{P}(\#\{1\leq j\leq N:|M_{ij}|\geq b_{N}^{k\epsilon}\}\geq 2NG(b_{N}^{k\epsilon}))\leq\exp(-\frac{NG(b_{N}^{k\epsilon})}{3})\leq\exp(-N^{\gamma})$
for $k\leq M$ and $\gamma=\frac{1}{4M+2}$. Moreover, applying this for every row and we have
(#$\mathbb{P}(\#\{1\leq j\leq N:|M_{ij}|\geq b_{N}^{k\epsilon}\}\geq 2NG(b_{N}^{k\epsilon})\text{ for some }i)\leq N\exp(-N^{\gamma}).$ (A.1)
Therefore,
$\sum_{j:|M_{i}j|\leq b_{N}^{\frac{M+1}{2M+1}}}|M_{ij}|\leq\sum_{k=0}^{M}\#(1\leq j\leq N:|M_{ij}|\geq b_{N}^{k\epsilon})b_{N}^{(k+1)\epsilon}$ (A.2)
$\leq\sum_{k=0}^{M}b_{N}^{(k+1)\epsilon}2NG(b_{N}^{k\epsilon})=\sum_{k=0}^{M}b_{N}^{(k+1)\epsilon}2NL(b_{N}^{k\epsilon})b_{N}^{-\alpha k\epsilon}\leq C_{1}b_{N}^{\epsilon}.$ (A.3)
This implies the first case. Case 2: $0<\alpha\leq 1$.
We again choose $\epsilon=\frac{1}{2M+1}$, where $\epsilon<\frac{\alpha}{8}$ and $\gamma=\frac{1}{4M+2}$. Since (A.1) also holds in this case,
$\sum_{j:|M_{ij}|\leq b_{N}^{\frac{M+1}{2M+1}}}|M_{ij}|\leq\sum_{k=0}^{M}\#(1\leq j\leq N:|M_{ij}|\geq b_{N}^{k\epsilon})b_{N}^{(k+1)\epsilon}$ (A.4)
$\leq\sum_{k=0}^{M}b_{N}^{(k+1)\epsilon}2NG(b_{N}^{k\epsilon})=\sum_{k=0}^{M}b_{N}^{(k+1)\epsilon}2NL(b_{N}^{k\epsilon})b_{N}^{-\alpha k\epsilon}\leq C_{2}b_{N}^{\epsilon}2Nb_{N}^{\frac{M(1-\alpha)}{2M+1}}\leq C_{2}b_{N}^{\frac{1}{2}+2\epsilon}.$ (A.5)
Proof of Lemma 3.5.
Given any $\epsilon>0$, we have the following:
$\mathbb{P}(\max_{1\leq i\leq j\leq N}|M_{ij}|<b_{N}^{1-\epsilon})=(1-\mathbb{P}(|M_{11}|>b_{N}^{1-\epsilon}))^{N(N+1)/2}$
$=O(\exp(-{\frac{N(N+1)}{2}\mathbb{P}(|M_{11}|>b_{N}^{1-\epsilon})})=O(\exp(-\frac{N(N+1)}{2}b_{N}^{(1-\epsilon)(-\alpha+\epsilon^{\prime})})))$
$=O(\exp(-N^{\delta}))$
for some $\delta>0$. This implies that
$\max_{1\leq i\leq j\leq N}|M_{ij}|>b_{N}^{1-\epsilon}$
holds with high probability for every $\epsilon>0$. Coupled with Lemma 3.4, this result leads to the following bound on the extreme eigenvalues of $M$:
$|\lambda_{N}|,\lambda_{1}\leq\|M\|_{\infty}\leq\max_{1\leq i\leq N}\sum_{1\leq j\leq N}|M_{ij}|=\max_{1\leq i\leq j\leq N}|M_{ij}|(1+O(b_{N}^{-1/8})).$
This proves the upper bounds of the desired lemma. For the lower bounds, we let $e_{k}$ be the $k$-th standard basis vector and let $f_{1}=\frac{1}{\sqrt{2}}(e_{i}+e_{j}),f_{2}=\frac{1}{\sqrt{2}}(e_{i}-e_{j})$. Then $|f_{1}^{*}Mf_{1}|,|f_{2}^{*}Mf_{2}|\geq\max|M_{ij}|(1+O(b_{N}^{-1/8}))$. Moreover, it is not hard to see that the signs of $f_{1}^{*}Mf_{1},f_{2}^{*}Mf_{2}$ are different. If $f_{1}^{*}Mf_{1}>0$, then $\lambda_{1}\geq f_{1}^{*}Mf_{1},\lambda_{n}\leq f_{2}^{*}Mf_{2}$. If $f_{1}^{*}Mf_{1}<0$,then $\lambda_{1}\geq f_{2}^{*}Mf_{2},\lambda_{n}\leq f_{1}^{*}Mf_{1}$. This and with high probability conditions about each rows in Lemma 3.3 conclude the proof of the lemma. ∎
Proof of Lemma 3.6.
Let
$M_{0}=\max_{1\leq i,j\leq N}{|M_{ij}|}$
$\mathbb{P}(\lambda_{1}<b_{N}N^{-\epsilon})$ $\leq\mathbb{P}(\lambda_{1}<b_{N}N^{-\epsilon},\lambda_{1}=M_{0}(1+O(b_{N}^{-1/8})))+\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))$
$\leq\mathbb{P}(M_{0}<2b_{N}N^{-\epsilon})+\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))$
Since Lemma 3.9 implies $b_{N}\geq N^{2/\alpha-\epsilon}$ for every $\epsilon>0$,
$\mathbb{P}(M_{0}<2b_{N}N^{-\epsilon})=(1-G(b_{N}N^{-\epsilon}))^{N(N+1)/2}=O(e^{-N^{\epsilon\alpha/2}})$
and due to Lemma 3.5
$\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))<N^{-\delta}$
for some $\delta>0$. These imply
$\mathbb{P}(\lambda_{1}<b_{N}N^{-\epsilon})<2N^{-\delta}$
and this implies $\lambda_{1}>b_{N}N^{-\epsilon}$ holds with high probability for every $\epsilon>0$.
For $x<y<x(1+Cx^{-\epsilon})$,
$(1-G(y))^{{N(N+1)}/{2}}-(1-G(x))^{N(N+1)/2}\leq(G(x)-G(y))\frac{N(N+1)}{2}(1-G(x))^{N(N+1)/2-1}$
$\ \ \ \leq\frac{G(x)-G(y)}{G(x)}\frac{N(N+1)}{2}G(x)(1-G(x))^{N(N+1)/2-1}$
Since $e^{x^{\delta}}L(x)$ is increasing function for $\delta=\epsilon/2$, $L(y)/L(x)\geq e^{x^{\delta}-y^{\delta}}\geq e^{\delta(x-y)x^{\delta-1}}$.
Hence,
$\frac{G(y)}{G(x)}\geq\exp(\alpha\log x-\alpha\log y+\delta(x-y)x^{\delta-1})\geq\exp(-C^{\prime}x^{-\epsilon/2}).$
This implies
$\frac{G(x)-G(y)}{G(x)}\leq 1-\exp(-C^{\prime}x^{-\epsilon/2})=O(x^{-\epsilon/2}).$
Also,
$\lim_{N\to\infty}\frac{N(N+1)}{2}G(x)(1-G(x))^{N(N+1)/2-1}=ue^{-u}$
where
$\lim_{N\to\infty}\frac{N(N+1)}{2}G(x)=u.$
Since $ue^{-u}$ is bounded, $\frac{N(N+1)}{2}G(x)(1-G(x))^{N(N+1)/2-1}$ is also bounded.
This means
$\mathbb{P}(x<M_{0}<y)=O(x^{-\epsilon/2})\text{ for }y=x(1+O(x^{-\epsilon})).$ (A.6)
For every $\epsilon>0$, since $\lambda_{1}>b_{N}N^{-\epsilon/2}$ holds with high probability, $\lambda_{1}-\lambda_{2}<b_{N}N^{-\epsilon}$ implies $\lambda_{2}>b_{N}N^{-\epsilon/2}-b_{N}N^{-\epsilon}$, this and (A.6) implies
$\mathbb{P}(\lambda_{1}-\lambda_{2}<b_{N}N^{-\epsilon})=\mathbb{P}(\lambda_{2}<\lambda_{1}<\lambda_{2}+b_{N}N^{-\epsilon})$
$\ \ \leq\mathbb{P}(\lambda_{2}<M_{0}(1+O(b_{N}^{-1/8}))<\lambda_{2}+b_{N}N^{-\epsilon})+\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))=O(N^{-\epsilon^{\prime}})$
for some $\epsilon^{\prime}>0$.
Similarly, (A.6) implies
$\mathbb{P}(Cb_{N}-b_{N}N^{-\epsilon}<\lambda_{1}<Cb_{N})$
$\leq\mathbb{P}(Cb_{N}-b_{N}N^{-\epsilon}<M_{0}(1+O(b_{N}^{-1/8}))<Cb_{N})+\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))$
$=O(N^{-\epsilon^{\prime}})$
for some $\epsilon^{\prime}>0$.
Proof of Lemma 3.9.
In order to prove the first part of the lemma, it suffices to show that for any $0<\epsilon<\alpha$ there exists $C_{\epsilon}>0$ such that $C_{\epsilon}t^{-\alpha-\epsilon}<\mathbb{P}(|X|>t)<C_{\epsilon}t^{-\alpha+\epsilon}$ holds for any $t>C_{\epsilon}$. For a given $\epsilon>0$, we can easily see from the definition of heavy-tailed random variable in Definition 2.2 that there exists $t_{0}>0$ such that
$\mathbb{P}(|X|>2t)/\mathbb{P}(|X|>t)<2^{-\alpha+\epsilon}$
holds for any $t>t_{0}$. For any $y>2t_{0}$ we can choose an integer $m$ such that $t_{0}<y/2^{m}\leq 2t_{0}$. Applying the inequality above $m$-times, we find that
$\mathbb{P}(|X|>y)/\mathbb{P}(|X|>y/2^{m})\leq 2^{m(-\alpha+\epsilon)}\leq(2t_{0}/y)^{\alpha-\epsilon}=y^{-\alpha+\epsilon}(2t_{0})^{\alpha-\epsilon}.$
Hence, we obtain
$\mathbb{P}(|X|>y)<\mathbb{P}(|X|>y/2^{m})y^{-\alpha+\epsilon}(2t_{0})^{\alpha-\epsilon}=\mathbb{P}(|X|>t_{0})(2t_{0})^{\alpha-\epsilon}y^{-\alpha+\epsilon}=C_{\epsilon}y^{-\alpha+\epsilon},$
which proves the desired upper bound. The proof for the lower bound is similar and we omit it here.
The second part of the lemma obviously holds due to the first part. This concludes the proof of the lemma. ∎
Proof of Lemma 3.10.
Let $Y=X\mathbbm{1}_{\{X\leq a_{N}\}}$, $Z=X\mathbbm{1}_{\{X>a_{N}\}}$ and let $f$ be the probability density function of X and $F(x)=\int_{x}^{\infty}f(s)\mathrm{d}s$
$\mathbb{E}(Y^{\delta})=\int_{0}^{a_{N}}x^{\delta}f\mathrm{d}x=[-x^{\delta}F(x)]_{0}^{a_{N}}+\int_{0}^{a_{N}}\delta x^{\delta-1}F(x)\mathrm{d}x$
For $\epsilon<\alpha-\delta$
$\int_{0}^{a_{n}}\delta x^{\delta-1}F(x)\mathrm{d}x\leq\int_{0}^{C_{\epsilon}}\delta x^{\delta-1}\mathrm{d}x+\int_{C_{\epsilon}}^{a_{n}}C_{\epsilon}\delta x^{\delta-1}x^{-\alpha+\epsilon}=O(1)$
where $C_{\epsilon}$ is such that for $\epsilon>0$ which satisfies $\mathbb{P}(X>t)\leq C_{\epsilon}t^{-\alpha+\epsilon}$ for every $t>C_{\epsilon}$
$\mathbb{E}(Z^{\delta})=\int_{a_{N}}^{\infty}x^{\delta}f\mathrm{d}x=[-x^{\delta}F]_{a_{N}}^{\infty}+\int_{a_{N}}^{\infty}\delta x^{\delta-1}F(x)\mathrm{d}x$
$\lim_{x\to\infty}x^{\delta}F(x)\leq\lim_{x\to\infty}C_{\epsilon}x^{\delta}x^{-\alpha+\epsilon}=0$ for $\epsilon<\alpha-\delta$ and
$\int_{a_{N}}^{\infty}\delta x^{\delta-1}F(x)\mathrm{d}x\leq\int_{a_{N}}^{\infty}\delta C_{\epsilon}x^{\delta-1}x^{-\alpha+\epsilon}\mathrm{d}x=O(a_{N}^{-\alpha+\delta+\epsilon}).$
Proof of Lemma 3.12.
Since
$\sum_{i}\lambda_{i}^{2}=\sum_{i,j=1}^{N}M_{ij}^{2}$
and $M_{ij}^{2}$ follows $\frac{\alpha}{2}$-stable laws and $\alpha/2<1$ holds,
for $\delta<\alpha/2<1$
$\mathbb{E}((\sum_{i}\lambda_{i}^{2})^{\delta})=O(N^{2})$
due to Lemma 3.11. Hence, for every $\epsilon>0$
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>N^{4/\alpha+\epsilon})N^{\delta(4/\alpha+\epsilon)}\leq\mathbb{E}((\sum_{i}\lambda_{i}^{2})^{\delta})=O(N^{2})$
Choose $\delta=\alpha/2-\epsilon/100$ and we have
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>N^{4/\alpha+\epsilon})\leq O(N^{-\alpha\epsilon/4}).$
Proof of Lemma 3.13.
Since
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>b_{N}^{2}N^{\epsilon})\leq O(N^{-\alpha\epsilon/8}),$
with high probability,
$\sum_{i}\lambda_{i}^{2}\leq b_{N}^{2}N^{\epsilon}$
Thus, with high probability,
#$\#\{|\lambda_{i}|>b_{N}N^{-\epsilon}\}=O(N^{3\epsilon}).$
Proof of Lemma 3.14.
For the sum of the square of eigenvalues, we have
$\sum_{i}\lambda_{i}^{2}=\operatorname{Tr}(MM^{*})=\sum_{i,j}a_{ij}^{2},$
where we set $a_{ij}=M_{ij}$. Let
$c_{N}=\inf\{t:\mathbb{P}(a_{ij}^{2}>t)<\frac{2}{N(N+1)}\}=b_{N}^{2}.$
By Proposition 3.7, for
$d_{N}=\frac{N(N+1)}{2}\mathbb{E}(a^{2}_{11}\mathbbm{1}_{a_{11}^{2}\leq c_{N}})=\frac{N(N+1)}{2}\mathbb{E}(a_{11}^{2}\mathbbm{1}_{|a_{11}|\leq b_{N}}).$
Since $a_{ij}^{2}$ follows $\alpha/2<1$ stable law, the remark implies
$d_{N}/c_{N}\to c$
for some constant c. Since
$\sum_{i,j=1}^{N}a_{ij}^{2}=2\sum_{1\leq i\leq j\leq N}a_{ij}^{2}-\sum_{i}a_{ii}^{2},$
this implies
$\frac{\sum_{i\leq j}a_{ij}^{2}-d_{N}}{c_{N}}\Rightarrow Y$
for some non-dengerate random variable Y.
Hence,
$\frac{\sum\lambda_{i}^{2}}{b_{N}^{2}}\Rightarrow X$
for some non-degenerate random variable X. ∎
Proof of Lemma 3.15.
Let us assume that each entry follows a $\alpha$- stable random variable X and $a_{N}=\inf\{u:1-\mathbb{P}(|X|>u)\leq 1/N\}$ and $c_{N}=N\mathbb{E}(X\mathbbm{1}_{|X|\leq a_{N}})$. The sum of eigenvalues satisfies
$\sum_{i=1}^{N}\lambda_{i}=\operatorname{Tr}(M)=\sum_{i=1}^{N}{M_{ii}}.$
For $\alpha<1$, Proposition 3.7 implies convergence of $\operatorname{Tr}(M)/a_{N}$ and $\lim_{N\to\infty}a_{N}/b_{N}=0$. This shows
$\lim_{N\to\infty}\sum_{i=1}^{N}\lambda_{i}/b_{N}=0.$
For $\alpha=1$,
$\mathbb{E}(X\mathbbm{1}_{|X|\leq a_{N}})=\int_{0}^{a_{N}}\mathbb{P}(|X|>u)\mathrm{d}u(1+o(1))=O(N^{\epsilon}).$
for every $\epsilon>0.$ Since $b_{N}\geq N^{2/\alpha-\epsilon}=N^{2-\epsilon}$ for every $\epsilon>0$, $\lim_{N\to\infty}c_{N}/b_{N}=0.$ This and convergence of $(\operatorname{Tr}(M)-c_{N})/a_{N}$ coming from Proposition 3.7 implies
$\lim_{N\to\infty}\operatorname{Tr}(M)/b_{N}=0$
For $\alpha>1$,
$\mathbb{E}(X\mathbbm{1}_{|X|\leq a_{N}})=\int_{0}^{a_{N}}\mathbb{P}(|X|>u)\mathrm{d}u(1+o(1))=O(1).$
Furthermore, $b_{N}\geq n^{2/\alpha-\epsilon}$ for every $\epsilon>0$ and $\alpha<2$ implies $\lim_{N\to\infty}N/b_{N}=0$ and $\lim_{N\to\infty}c_{N}/b_{N}=0.$ Therefore, convergence of $(\operatorname{Tr}(M)-c_{N})/a_{N}$ coming from Proposition 3.7 and the above implies
$\lim_{n\to\infty}\operatorname{Tr}(M)/b_{N}=0.$
Proof of Lemma 3.16.
We can diagonalize the interaction matrix $M=O^{T}DO$ for an orthogonal matrix O and a diagonal matrix $D=\operatorname{diag}(\lambda_{1},..,\lambda_{N})$ and apply the change of variable $x\to O^{-1}x$ then we have
$Z_{n}$ $=\frac{1}{|S_{N}|}\int e^{\beta b_{N}^{-1}<x,Mx>}\mathrm{d}\Omega=\frac{1}{|S_{N}|}\int_{S_{N}}e^{\beta b_{N}^{-1}\sum_{i}\lambda_{i}x_{i}^{2}}.$
We will use Laplace transform to calculate above statistics, so we define below statistics.
$J(z)$ $=\int_{\mathbb{R}^{N}}e^{\beta N\sum_{i}\lambda_{i}y_{i}^{2}}e^{-\beta Nz\sum_{i}y_{i}^{2}}\mathrm{d}y,\ \operatorname{\mathrm{Re}}z>\lambda_{1}$
$I(t)$ $=\int_{\mathbb{S}^{N-1}}e^{t\sum_{i}\lambda_{i}x_{i}^{2}}\mathrm{d}\Omega$
$J(z)$ is calculated as
$J(z)$ $=\int_{\mathbb{R}^{N}}e^{\beta N\sum_{i}(\lambda_{i}-z)y_{i}^{2}}\mathrm{d}y=(\frac{\pi}{\beta N})^{N/2}\prod_{i}\frac{1}{\sqrt{z-\lambda_{i}}},\operatorname{\mathrm{Re}}z>\lambda_{1}$
and
[…]
Now we apply Laplace transform and we have
$\frac{t^{N/2-1}I(t)}{2(\beta N)^{N/2}}$ $=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}e^{zt}J(z)\mathrm{d}z=\left(\frac{\pi}{\beta N}\right)^{N/2}\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}e^{zt}\prod\frac{1}{\sqrt{z-\lambda_{i}}}\mathrm{d}z.$
On the other hand,
$Z_{N}$ $=\frac{1}{S^{N-1}}\int_{S^{N-1}}e^{b_{N}^{-1}\beta N\sum_{i}\lambda_{i}x_{i}^{2}}\mathrm{d}\sigma=\frac{1}{|S^{N-1}|}I(\frac{\beta N}{b_{N}})$
$=\frac{\Gamma(N/2)}{2\pi^{N/2}}\cdot\frac{2\pi^{N/2}b_{N}^{N/2-1}}{(\beta N)^{N/2-1}}\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}e^{\frac{\beta Nz}{b_{N}}}\prod_{i=1}^{N}\frac{1}{\sqrt{z-\lambda_{i}}}\mathrm{d}z$
$=C_{N}\frac{1}{i}\int_{\gamma-i\infty}^{\gamma+i\infty}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z$
where
$G(z)=2\beta z-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(z-\lambda_{i}),\ C_{N}=\frac{\Gamma(N/2)b_{N}^{N/2-1}}{2\pi(N\beta)^{N/2-1}}.$
Proof of Lemma 3.17.
We first consider the first case, $\lambda_{1}<\frac{b_{N}}{2\beta}$. From Lemma 3.13,
$0\leq\sum_{|\lambda_{i}|>b_{N}N^{-\epsilon}}\frac{1}{z-\lambda_{i}}\leq\#\{|\lambda_{i}|>b_{N}N^{-\epsilon}\}\frac{1}{z-\lambda_{1}}=O(\frac{N^{3\epsilon}}{z-\lambda_{1}}).$
Since we have
$z>\lambda_{1}>b_{N}N^{-\epsilon/4}$
with high probability due to Lemma 3.6,
$\sum_{|\lambda_{i}|<b_{N}N^{-\epsilon}}\frac{1}{z-\lambda_{i}}=(N-O(N^{3\epsilon}))\frac{1}{z}(1+O(N^{-3/4\epsilon}))=\frac{N(1+O(N^{-3/4\epsilon}))}{z}.$
Summing up them and we have
$G^{\prime}(z)=2\beta-\frac{b_{N}}{N}(O(\frac{N^{3\epsilon}}{z-\lambda_{1}})+\frac{N(1+O(N^{-3/4\epsilon})}{z})$
$G^{\prime}(\frac{b_{N}}{2\beta}-b_{N}N^{-\epsilon/2})<0,\ G(\frac{b_{N}}{2\beta}+b_{N}N^{-\epsilon/2})>0$
Here $\lambda_{1}<\frac{b_{N}}{2\beta}$ implies $\lambda_{1}<\frac{b_{N}}{2\beta}-b_{N}N^{-\epsilon/4}$ holds with high probability due to Lemma 3.6. Since $\lambda_{1}=\max\{|M_{ij}|(1+O(b_{N}^{-1/8}))\}$ and $|\lambda_{N}|=\max\{|M_{ij}|(1+O(b_{N}^{-1/8}))\}$ due to Lemma 3.5,
$\gamma=\frac{b_{N}}{2\beta}(1+O(N^{-\epsilon/2}))>\lambda_{1},|\lambda_{N}|$
holds with high probability. Hence,
$G^{\prime}(\gamma)=2\beta-\frac{b_{N}}{N}\sum_{i=1}^{N}\frac{1}{\gamma-\lambda_{i}}=2\beta-\frac{b_{N}}{N\gamma}(N+\sum_{i=1}^{N}\sum_{k=1}^{\infty}\lambda_{i}^{k}/\gamma^{k}).$
Let us show that
$X_{N}=\sum_{i=1}^{N}\sum_{k=1}^{\infty}\lambda_{i}^{k}/\gamma^{k}$
converges to non degenerate random variable X.
Lemma 3.15 implies
$\lim_{N\to\infty}\sum_{i=1}^{N}\lambda_{i}/\gamma=0.$
The remaining terms are
$\left|\frac{\sum_{i=1}^{N}\lambda_{i}^{k}}{\gamma^{k}}\right|=\left|\frac{\sum_{i=1}^{N}\lambda_{i}^{k}}{(b_{N}/2\beta)^{k}}(1+O(N^{-\epsilon}))\right|\leq\frac{\Gamma^{k-2}\sum_{i=1}^{N}\lambda_{i}^{2}}{(b_{N}/2\beta)^{k}}(1+O(N^{-\epsilon}))$
where $\Gamma=\max\{|\lambda_{1}|,|\lambda_{N}|\}$.
Since $\Gamma/(b_{N}/2\beta)$ converges to same random variable with convergence of $\lambda_{1}/(b_{N}/2\beta)$ and Lemma 3.14 implies $\sum_{i=1}^{N}\lambda_{i}^{2}/(b_{N}/2\beta)^{2}$ converges to a random variable,
$|X_{N}|\leq\sum_{i}\lambda_{i}/\gamma+\frac{\Gamma/(b_{N}/2\beta)}{1-(\lambda_{1}/(b_{N}/2\beta))}\sum_{i=1}^{N}(\lambda_{i}/(b_{n}/2\beta))^{2}(1+O(N^{-\epsilon}))$
implies that its upper bound has convergence random variable.
This implies
$0=G^{\prime}(\gamma)=2\beta-\frac{b_{N}}{N\gamma}(N+X_{N})$
and $X_{N}$ converges to a non degenerate random variable X.
Thus,
$\gamma=\frac{b_{N}}{2\beta}+\frac{b_{N}X_{N}}{2\beta N}$
Case 2: $\lambda_{1}>\frac{b_{N}}{2\beta}$ holds.
$\sum_{|\lambda_{i}|>b_{N}N^{-\epsilon}}\frac{1}{z-\lambda_{i}}=\frac{1}{z-\lambda_{1}}+\frac{O(N^{3\epsilon})}{z-\lambda_{2}}$
Summing up and we obtain
$G^{\prime}(z)=2\beta-\frac{b_{N}}{N}(\frac{1}{z-\lambda_{1}}+O(\frac{N^{3\epsilon}}{z-\lambda_{2}})+\frac{N(1+O(N^{-3/4\epsilon})}{z})$
Since $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$ holds with high probability due to Lemma 3.6,
$G^{\prime}(z)=2\beta-\frac{b_{N}}{N}(\frac{1}{z-\lambda_{1}}+O(\frac{N^{4\epsilon}}{b_{N}})+\frac{N(1+O(N^{-3/4\epsilon})}{z})$
for $z>\lambda_{1}$. Hence,
$G^{\prime}(\lambda_{1}+\frac{1}{2\beta-\frac{b_{N}}{\lambda_{1}}}\frac{b_{N}}{N}+\frac{b_{N}}{N}N^{-\epsilon/2})>0,G^{\prime}(\lambda_{1}+\frac{1}{2\beta-\frac{b_{N}}{\lambda_{1}}}\frac{b_{N}}{N}-\frac{b_{N}}{N}N^{-\epsilon/2})<0.$
This implies
$\gamma=\lambda_{1}+\frac{1}{2\beta-\frac{b_{N}}{\lambda_{1}}}\frac{b_{N}}{N}+O(\frac{b_{N}}{N}N^{-\epsilon/2}).$
Appendix B Statistical property for T
We also want to know the property of the statistics T. Let $T_{N}=-\sum_{i}\log(1-\frac{\lambda_{i}}{\gamma})$
$\mathbb{E}(-\log(1-\sum_{i=1}^{N}\lambda_{i}/\gamma))=\sum_{k=1}^{\infty}\frac{1}{k}\mathbb{E}(\sum_{i=1}^{N}(\lambda_{i}/\gamma)^{k})$
In this case is the conditional case and every entry of interaction matrix follows the $Y=X1_{|X|\leq\frac{b_{N}}{2\beta}}$.
$\lim_{N\to\infty}\mathbb{E}((\frac{b_{N}}{2\beta})^{-k}\sum\lambda_{i}^{k})=\begin{cases}0&\textit{k is odd}\\ \frac{k}{k-\alpha}&\textit{k is even }\end{cases}$
$\mathbb{E}(\sum_{i=1}^{N}\lambda_{i}^{k})=\mathbb{E}(TrM^{k})=\mathbb{E}(\sum M_{i_{1}i_{2}}...M_{i_{k}i_{1}})$
Since k is finite, if $\{i_{1},...,i_{k}\}$ has t vertices, then the possible choice is $n(n-1)...(n-t+1)$.
[…]
For $k\geq 2$
$\mathbb{E}(|Y|^{k})=\int_{0}^{b_{N}/2\beta}ku^{k-1}\mathbb{P}(|Y|>u)\mathrm{d}u=\frac{(2\beta)^{\alpha}}{k-\alpha}(\frac{b_{N}}{2\beta})^{k}\frac{2k}{N(N+1)}(1+o(1))$
Therefore for the corresponding graph $T$,
$\mathbb{E}(\sum_{T}M_{i_{1}i_{2}}...M_{i_{k}i_{1}})=\prod_{x=1}^{t}(N-x+1)\mathbb{E}(e_{1}^{t_{1}}...e_{l}^{t_{l}})=\prod_{x=1}^{t}(N-x+1)\prod_{i=1}^{l}\mathbb{E}(Y^{t_{i}})$
For $\alpha<1$,
$\mathbb{E}(\sum_{T}M_{i_{1}i_{2}}...M_{i_{k}i_{1}})=(\frac{b_{N}}{2\beta})^{k}\prod_{x=1}^{t}(N-x+1)\prod_{i=1}^{l}\frac{2(2\beta)^{\alpha}t_{i}}{(t_{i}-\alpha)N(N+1)}$
$\lim_{N\to\infty}\mathbb{E}((\frac{b_{N}}{2\beta})^{-k}\sum_{T}M_{i_{1}i_{2}}...M_{i_{k}i_{1}})$
it converges not to 0, only if $t=2l$. $t$ is the number of vertices, and $l$ is the distinct kind of edges in $\{i_{1}i_{2},...,i_{k}i_{1}\}$.Since the number of edges are at least $t-1$, $t=2l\geq 2t-2$ and $t\leq 2$. Hence, the only possible case is $l=1,t=2.$ This case is only possible when k is even and when the graph T is $k$ edges between two vertices. Combining altogether, we have
$\lim_{N\to\infty}\mathbb{E}((\frac{b_{N}}{2\beta})^{-k}(\sum_{i=1}^{N}\lambda_{i}^{k}))=\begin{cases}0&k\textit{ is odd}\\ \frac{2(2\beta)^{\alpha}k}{k-\alpha}&k\textit{ is even}\end{cases}$
Therefore,
$\lim_{N\to\infty}\mathbb{E}(-\sum_{i=1}^{N}\log(1-\frac{\lambda_{i}}{b_{N}/2\beta}))=\lim_{N\to\infty}\mathbb{E}(\sum_{k=1}^{\infty}\frac{1}{k}\sum(\lambda_{i}/(b_{N}/2\beta))^{k})=\sum_{t=1}^{\infty}\frac{2(2\beta)^{\alpha}}{(2t-\alpha)}$
Similarly for $\alpha\geq 1$, let $l^{\prime}$ be the number of edges appear once in $T$.
$\mathbb{E}(\sum_{T}M_{i_{1}i_{2}}...M_{i_{k}i_{1}})=(\frac{b_{N}}{2\beta})^{k-l^{\prime}}\prod_{x=1}^{t}(N-x+1)\prod_{i=1}^{l-l^{\prime}}\frac{2(2\beta)^{\alpha}t_{i}}{(t_{i}-\alpha)N(N+1)}O(1)^{l^{\prime}}$
To make the limit
$\lim_{N\to\infty}\mathbb{E}((\frac{b_{N}}{2\beta})^{-k}\sum_{T}M_{i_{1}i_{2}}...M_{i_{k}i_{1}})$
not to be 0, it should satisfy $l+1\geq t\geq l^{\prime}(2/\alpha)+2(l-l^{\prime})$. Let $l^{\prime\prime}=l-l^{\prime}$, then $l^{\prime}+l^{\prime\prime}+1\geq(2/\alpha)l^{\prime}+2l^{\prime\prime}$. Since $l^{\prime\prime}+(\alpha/2-1)\leq 1$, $l^{\prime\prime}=0\textit{ or }1$. If $l^{\prime\prime}=0$, then
$\lim_{N\to\infty}\mathbb{E}((\frac{b_{N}}{2\beta})^{-k}\sum_{T}M_{i_{1}i_{2}}...M_{i_{k}i_{1}})=(\frac{b_{N}}{2\beta})^{-k}\prod_{x=1}^{t}(N-x+1)O(1)^{k}=0$
since $t\leq k<\frac{2}{\alpha}k$. Therefore, the remaining part is $l^{\prime\prime}=1,l^{\prime}=0$. This is same with above. Now we have
$\lim_{N\to\infty}\mathbb{E}(-\sum_{i=1}^{N}\log(1-\frac{\lambda_{i}}{b_{N}/2\beta}))=\lim_{N\to\infty}\mathbb{E}(\sum_{k=1}^{\infty}\frac{1}{k}\sum_{i=1}^{N}(\lambda_{i}/(b_{N}/2\beta))^{k})=\sum_{t=1}^{\infty}\frac{2(2\beta)^{\alpha}}{(2t-\alpha)}$
Due to Lemma 3.13, $\lim_{N\to\infty}\sum_{i=1}^{N}\{\log(1-\frac{\lambda_{i}}{b_{N}/2\beta})-\log(1-\frac{\lambda_{i}}{\gamma})\}=0.$ This implies
$\mathbb{E}(T)=\lim_{N\to\infty}\mathbb{E}(\frac{1}{2}T_{N})=\sum_{t=1}^{\infty}\frac{(2\beta)^{\alpha}}{(2t-\alpha)}$
for any cases. Now let us get back to calculate the limit of $\mathbb{E}(T^{2})$. To calculate this, we calculate
$\mathbb{E}(\sum_{i=1}^{N}\lambda_{i}^{k}\sum_{i=1}^{N}\lambda_{i}^{l})=\mathbb{E}(\sum_{T_{1}\cup T_{2}}(M_{i_{1}i_{2}}...M_{i_{k}i_{1}})(M_{j_{1}j_{2}}...M_{j_{l}j_{1}}))$
Let t be the number of elements in $\{i_{1}i_{2},...,i_{k-1}i_{k},i_{k}i_{1},j_{1}j_{2},..,j_{l-1}j_{l},j_{l}j_{1}\}$ and let s be the number of different edges. For fixed $T_{1}\cup T_{2}$, we also have similar equation. For $\alpha<1$
$\mathbb{E}((\frac{b_{N}}{2\beta})^{-k-l}\sum_{T_{1}\cup T_{2}}(M_{i_{1}i_{2}}...M_{i_{k}i_{1}})(M_{j_{1}j_{2}}...M_{j_{l}j_{1}})))=\prod_{x=1}^{t}(N-x+1)\prod_{i=1}^{s}\frac{2(2\beta)^{\alpha}t_{i}}{(t_{i}-\alpha)N(N+1)}$
It converges only when t=2s, and $s\geq t-2=2s-2$. The possible cases to above equation converges are $(t,s)=(2,1),(4,2)$. More precisely, the possible cases are
  • $T_{1}$ consists of $k$ edges $e_{1}$, $T_{2}$ consists of $l$ edges $e_{2}\neq e_{1}$.
  • $T_{1}$ consists of $k$ edges $e$.
$k,l$ are even number for both cases. Therefore,
[…]
For $\alpha\geq 1$ case, for the similar reason from the calculation of expectation, it also has the same value as above.
Since
$(\sum\log(1-\lambda_{i}/(b_{N}/2\beta))^{2}=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}(\frac{b_{N}}{2\beta})^{-k-l}\frac{1}{kl}\sum\lambda_{i}^{k}\sum\lambda_{i}^{l},$
$\mathbb{E}(T^{2})$ $=\lim_{N\to\infty}\mathbb{E}(\frac{1}{4}T_{N}^{2})=\lim_{N\to\infty}(\sum\log(1-\lambda_{i}/(b_{N}/2\beta))^{2}$
$=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\frac{((2\beta)^{\alpha})^{2}4kl}{2k(2k-\alpha)2l(2l-\alpha)}+\frac{(2\beta)^{\alpha}(k+l)}{4kl(2k+2l-\alpha)}.$
Moreover, the variance is
$V(T)=\lim_{N\to\infty}V(\frac{1}{2}T_{N})=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\frac{(2\beta)^{\alpha}(k+l)}{4kl(2k+2l-\alpha)}=\sum_{n=1}^{\infty}\frac{(2\beta)^{\alpha}}{2(2n+2-\alpha)}\sum_{i=1}^{n}\frac{1}{i}.$
Acknowledgement
The first author thanks to Ji Hyung Jung for helpful discussions. The work was partially supported by National Research Foundation of Korea under grant number NRF-2019R1A5A1028324 and NRF-2023R1A2C1005843.